OpenStudy (anonymous):

a child lies on the ground and looks up at the top o a 17ft tree nearby. the child is 14ft away from the child to the top of the tree? round to the nearest whole degree.

OpenStudy (anonymous):

theres some part missing i think?

OpenStudy (anonymous):

no thats all

OpenStudy (anonymous):

i guess it is the child is 14ft away from the base of the tree. find the distance from the child to top of tree. It yes , then the answer is distance between child and top of tree will be the hypotenuse of a right angles triangle whose altitude/perpendicular is 17ft and base is 14 ft then H^2 = 17^2 + 14^2 = 289 + 196 = 485 so H = root485 = 22.02 ft.

OpenStudy (anonymous):

but he asks to round to the nearest degree, so i suppose he asks for an angle... are you sure that there isnt a part missing between "the child is 14ft away from" and "the child to the top of the tree"?

OpenStudy (anonymous):

yeah, i missed out that....sorry